OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2sin^2x = sin x I'm soo lost..

OpenStudy (anonymous):

what do i do after i get to pi/6?

hartnn (hartnn):

2sin^2x - sin x = 0 factor out sin x

hartnn (hartnn):

sin x (2sin x-1) = 0 sin x = 0 for which angles of x ? sin x = 1/2 for which angles of x ?

OpenStudy (anonymous):

.....?

OpenStudy (anonymous):

i said sinx = 1/2 at pi/6

OpenStudy (anonymous):

now what do i do? i don't understand what it means to = 0? i didn't think it did... idk

hartnn (hartnn):

sin x can be 0 too, right ? so, x can take values of 0 and pi too, isn't it ?

hartnn (hartnn):

also, there are 2 values of x for which sin x = 1/2 pi/6 is one of them, what is other ?

OpenStudy (anonymous):

uhm... 5pi/6

OpenStudy (anonymous):

and just pi is an answer too?

hartnn (hartnn):

5pi/6, yes! so your 4 answers are 0, pi/6, pi, 5pi/6

hartnn (hartnn):

didn't you get why sin x =0 can also be true ?

OpenStudy (anonymous):

no i got all the other answers i just dont understand why it can equal 0

hartnn (hartnn):

2sin^2x - sin x = 0 factor out sin x sin x(2 sin x -1) = 0 when a b = 0 a=0 or b= 0 so, when sin x(2 sin x -1) = 0 sin x = 0 or 2 sin x -1 = 0 and since sin x = 0, we get x = 0 , pi

OpenStudy (anonymous):

ohhh okay thank you. i think i get it.